Liquidation

Southwestern Wear Inc. has the following balance sheet:

Current assets $1,875,000 Accounts payable $375,000
Fixed assets 1,875,000 Notes payable 750,000
Subordinated debentures 750,000
Total debt $1,875,000
Common equity 1,875,000
Total assets $3,750,000 Total liabilities and equity $3,750,000

The trustee's costs total $331,750, and the firm has no accrued taxes or wages. Southwestern has no unfunded pension liabilities. The debentures are subordinated only to the notes payable. If the firm goes bankrupt and liquidates, how much will each class of investors receive if a total of $3 million is received from sale of the assets?

Distribution of proceeds on liquidation:

1. Proceeds from sale of assets $
2. First mortgage, paid from sale of assets $
3. Fees and expenses of administration of bankruptcy $
4. Wages due workers earned within 3 months
prior to filing of bankruptcy petition
$
5. Taxes $
6. Unfunded pension liabilities $
7. Available to general creditors $

Distribution to general creditors:

Claims of General Creditors
Claim
(1)
Application of 100% Distribution
(2)
After Subordination Adjustment
(3)
Percentage of Original Claims Received
(4)
Notes payable $ $ $ %
Accounts payable $ $ $ %
Subordinated debentures $ $ $ %
Total $ $ $

The remaining $ will go to the common stockholders.
Answer #1

1)3000000
2)0
3)331750
4)0
5)0
6)0
7)=3000000-331750=2668250
Of the total amount the preference is below:
Accounts payable=375000
notes payable=750000
Subordinated debenture=75000
the remaining left will go to the common shareholders and it is =2668250-375000-750000-750000
=793250

Claims of General Creditors Notes payable Accounts payable Subordinated debentures Tota Application of 100% After Subordination Percentage of Original Claim(1) Distribution(2 Adjustment(3 Claims Received(4 750,000 375,000 750,000 1,875,000 28.11% 14.05% 28.11% 7027% 750,000 375,000 750,000 1875000 25.00% 12.50% 25.00% 62.50% Claims of General Creditors Claim(1) 750000 375000 4 Subordinated debentures 750000 Application of 100% Distribution(2) After Subordination Adjustment(3) Percentage of Original Claims Received(4) 750000 375000 750000 SUM(D2:D4) 2 Notes payable 3 Accounts payable B2/2668250 -B3/2668250 B426682 50 SUM(B2:84) SUM(C2:C4) D2/3000000 D3/3000000 D4/3000000 SUM E2:E4) :5 Total